OpenStudy (anonymous):

Which is a zero of the absolute value function f(x) = 1 - 3|3x5 - 4x2|?

OpenStudy (paxpolaris):

\[\large 0=1-3\left| 3x^5-4x^2 \right|\]\[\large \implies 3x^2\left( 3x^3-4 \right)=\pm1\] So, solve, \[\large 9x^5-12x^2-1=0\ and\ \large 9x^5-12x^2+1=0\]

OpenStudy (anonymous):

solve if i solve this i will get my answer

OpenStudy (paxpolaris):

yes ... *and should be or